On July 1, 20X2, Alan Enterprises merged with Terry Corporation through an exchange of stock and the subsequent liquidation of Terry. Alan issued 200,000 shares of its stock to effect the combination. The book values of Terry’s assets and liabilities were equal to their fair values at the date of combination, and the value of the shares exchanged was equal to Cherry’s book value. Information relating to income for the companies is as follows:

20X1 Jan. 1–June 30, 20X2 July 1–Dec. 31, 20X2
Net Income:
Alan Enterprises $ 4,460,000 $ 2,500,000 $ 3,528,000
Terry Corporation 1,300,000 692,000


Alan Enterprises had 1,000,000 shares of stock outstanding prior to the combination. Remember that when calculating earnings per share (EPS) for the year of the combination, the shares issued in the combination were not outstanding for the entire year.

Required:
Compute the net income and earnings-per-share amounts that would be reported in Alan's 20X2 comparative income statements for both 20X2 and 20X1. (Round earnings per share to 2 decimal places.)
Answer #1

ANSWER

Computation of Net Income and EPS for 20X2
Net Income= 2,500,000+3,528,000= $6,028,000
EPS= Net Income / No. of Share
=6,028,000/(1,000,000 +200,000 *12/6)
=$4.30
Computation of Net Income and EPS for 20X1
Net Income= 4,460,000
EPS= Net Income / No. of Share
=4,460,000/1,000,000
=$4.46

__________________________________

If you have any query or any Explanation please ask me in the comment box, i am here to helps you.please give me positive rating.

*****************THANK YOU*************
Answer #2

The previous answer is almost correct. EPS for 20X2 is more accurately measured by dividing the NI for that year by 1 million shares plus half the newly issued shares, according to the solutions manual. 


EPS 20X2= 6028000/(1000000+100000)=$5.48